OpenStudy (anonymous):

Solve rational equations \frac(9)(x+1)-\frac(5)(2) = \frac(2)(3x+3).

OpenStudy (fibonaccichick666):

\[\frac{9}{x+1}-\frac{5}{2} = \frac{2}{3x+3} \]

OpenStudy (fibonaccichick666):

yea?

OpenStudy (anonymous):

yes

OpenStudy (fibonaccichick666):

btw you were close with the LAtex, you need to use curly brackets instead of parentheses and add a \bracket to both ends

OpenStudy (fibonaccichick666):

alright, so anyways, what have you tried so far?

OpenStudy (anonymous):

well im confused since theres two x.

OpenStudy (anonymous):

i know from 3x+3 its 3(x+1)

OpenStudy (fibonaccichick666):

that is true

OpenStudy (fibonaccichick666):

so you have a few ways to go about it

OpenStudy (anonymous):

do i multiply by 6(x+1)

OpenStudy (fibonaccichick666):

well you could, but you wont really need the 6, it does eliminate all of the fractions though

OpenStudy (fibonaccichick666):

I mean it is legal and it works

OpenStudy (fibonaccichick666):

let me know what you get when you do

OpenStudy (anonymous):

help me please with number 1

OpenStudy (fibonaccichick666):

(don't forget to use parentheses so you hit every term!)

OpenStudy (anonymous):